In 2019, Rory earns $74,000 per year as a college professor, paying $6,771 in federal income taxes. Latesia is a marketing executive with a salary of $144,000 who paid $22,879 in federal taxes. In addition, Latesia collected $100,000 on a life insurance policy. Both taxpayers are single and take the 2019 standard deduction.

Compute the following for the two taxpayers.

Note: Round the total taxes to two decimal places. Enter percentages to one decimal.

a. With respect to the Social Security tax, calculate their total FICA liability.

Hint: OASDI: 6.2% with a cap of $132,900; MHI: 1.45%

Rory's total FICA taxes: $
Latesia's total FICA taxes: $

Click here to view the 2019 Tax Rate Schedules (for single taxpayers).

b. Based on the data provided, compute Rory and Latesia's marginal tax rate for 2019.

Rory's marginal tax rate: %
Latesia's marginal tax rate: %

c. Based on the data provided, compute Rory and Latesia's average tax rate for 2019.

Rory's average tax rate: %
Latesia's average tax rate: %

d. Based on the data provided, compute Rory and Latesia's effective tax rate for 2019.

Rory's effective tax rate: %
Latesia's effective tax rate:
Answer #1
a. Total FICA liability.
Hint: OASDI: 6.2% with a cap of $132,900; MHI: 1.45%
Rory's total FICA taxes:(74000*(6.2%+1.45%)) 5661
Latesia's total FICA taxes:(132900*6.2%)+(144000+100000)*1.45% 11777.80
b. Based on the data provided, Rory and Latesia's marginal tax rate for 2019.
Rory's marginal tax rate:Taxable income =74000-std. deduction for single tax payers, 12200=61800 falls in 22% tax bracket, ie. Taxable income > $ 39475 < $ 84200
Latesia's marginal tax rate:(144000+100000)=244000-12200=231800 falls in 35% tax bracket, ie. Taxable income > $ 204100 < $ 510300
c. Based on the data provided, Rory and Latesia's average tax rate for 2019.
Rory's average tax rate: Av. Of diff. tax brackets ,ie. (10%+12%+22%)/3= 14.7%
Latesia's average tax rate:(10%+12%+22%+24%+32%+35%)/6= 22.5%
d. Based on the data provided, Rory and Latesia's effective tax rate for 2019.
Rory's effective tax rate:Actual tax expense/Total income,ie.6771/74000= 9.2%
Latesia's effective tax rate:22879/244000= 9.4%